One three consecutive passes, a football team gains 6 yards, loses 34 yards and gains 11 yards. What number represents the total net yardage?
Found 2 solutions by josmiceli, edjones:
Answer by josmiceli(19441)

+6 -34 + 11 = -34 + 17
and
-34 + 17 = -17

Answer by edjones(8007)

6-34+11=-17 net yardage